OpenStudy (anonymous):

during a basketball shooting contest, each contestant gets 10 shots. for each shot he attempts, parik's probability of making a basket is 55%. What is the probability he will make at least 8 shots in his 10 attempts? express the answer as a percent. A.) 0.3% B.) 1.6% C.) 8.4% D.) 10.0%

OpenStudy (anonymous):

at least 8 shots means you have have to compute the probability he makes 8, 9 and 10. We are assuming the shots are independent so use binomial. \[P(x=8)=\dbinom{10}{8}(.55)^8(.45)^2\]

OpenStudy (anonymous):

i can write out the others if you like.

OpenStudy (anonymous):

OKAY PLEASE :)

OpenStudy (anonymous):

the general formula is \[P(x=k)=\dbinom{n}{k}p^k(1-p)^{n-k}\] i used it for n = 10, k = 8 and p=.55 for k = 9 you get \[P(x=9)=\dbinom{10}{9}(.55)^9(.44)^{1}\]

OpenStudy (anonymous):

and k= 19 you get \[\dbinom{10}{10}(.55)^{10}\]

OpenStudy (anonymous):

\[\dbinom{10}{8} = \frac{10\times 9}{2}=45\] \[\dbinom{10}{9} = 10\] \[\dbinom{10}{10}=1\] the rest you need a calculator for.

OpenStudy (anonymous):

OKAY so what do i put in the caculator? im so confused

OpenStudy (anonymous):

oh sorry . you need to put the \[(.55)^8(.45)^2\] part in the calculator because you certainly cannot do it by hand. answer is \[45 \times (.55)^8(.45)^2 + 10 \times (.55)^9(.45)+(.55)^{10}\]

OpenStudy (anonymous):

i got .0995 rounded to .10= 10%

OpenStudy (anonymous):

thanx so much u deserve a medal

OpenStudy (anonymous):

you are welcome. thank you. btw if it is not clear why this formula is the correct one let me know. it is not too hard to explain.

OpenStudy (anonymous):

i just have one more problem to do and im not to good with probability Karin is playing a game at an after school carnival. There are a number of table tennis balls in a bag, and 1/3 of them are marked as prize-winners. A player randomly selects a ball from the bag, checks to see if they won a prize, and then replaces the ball. If karin plays the game 12 times, what is the probability that she will win a prize exactly 4 times? A.) 0.0004 B.) 0.012 C.) 0.238 D.)0.25